OpenStudy (anonymous):

help Find the geometric mean of 162 and 8. answears are 36, 56, 41, or 46 i never heard of this in my whole life so lol help plz.

Parth (parthkohli):

You basically set up a proportion here. \(\Large \color{MidnightBlue}{\rightarrow 8:x::x:162 }\) \(\Large \color{MidnightBlue}{\rightarrow x^2 = 162*8 }\) \(\Large \color{MidnightBlue}{\rightarrow x^2 = 1296 }\) \(\Large \color{MidnightBlue}{\rightarrow x = \sqrt{1296} }\) \(\Large \color{MidnightBlue}{\rightarrow x = 36 }\)

OpenStudy (callisto):

geometric mean \[=\sqrt{162 \times 8} = \sqrt{1296} = 36\]

Parth (parthkohli):

Always, a common ratio is in proportion to two numbers.

OpenStudy (anonymous):

lol wow that was easy XD

Parth (parthkohli):

:P